5. A firm is considering Projects S and L, whose cash flows are shown below. These projects are mutually exclusive, equally risky, and not repeatable. The CEO wants to use the IRR criterion, while the CFO favors the NPV method. You were hired to advise the firm on the best procedure. If the wrong decision criterion is used, how much potential value would the firm lose?
WACC: 6.00%
Year 0 1 2 3 4

CFS -$1,025 $380 $380 $380 $380

CFL -$2,150 $765 $765 $765 $765

a. $188.68
b. $198.61
c. $209.07
d. $219.52
e. $230.49
1
Expert's answer
2012-02-07T15:48:43-0500
c. $209.07

NPV:
PV[S] = -1025+380•∑(1/1.06^n) ≈ +291.74
PV[L] = -2150+765•∑(1/1.06^n) ≈ +500.8
PV[L] > PV[S]

IRR:
0 = -1025+380•∑(1/(1+IRR)^n) → IRR[S]≈17.861%
0 = -2150+765•∑(1/(1+IRR)^n) → IRR[L]≈15.781%
IRR[L] < IRR[S]
15.781 < 17.861

Since projects are mutually exclusive - then project [L] is preferred (despite lower
IRR - higher investment volume pays out), obviously assuming firm doesn't have any other investment opportunities.

Δ ≈ 500.8 - 291.74 = 209.07